Rigid rods of negligible mass lying along the y ax
Answer #1

a) I = I1 + I2 + I3

= 6.2*4.65^2 + 3.1*3.1^2 + 4.65*6.2^2

= 342.6 kg.m^2

b) KE(rotational) = 0.5*I*w^2

= 0.5*342.6*4^2

= 2740.8 J


c)

v1 = 4.65*4 = 18.6 m/s

v2 = 3.1*4 = 12.4 m/s

v3 = 6.2*4 = 24.8 m/s


d) total KE = 0.5*m1*v1^2 + 0.5*m2*v2^2 + 0.5*m3*v3^2

= 0.5*6.2*18.6^2 + 0.5*3.1*12.4^2 + 0.5*4.65*24.8^2

= 2740.8 J

e) 3rd option is corrrect answer

net 0.5*m*vi^2 = 0.5*I*w^2